OpenStudy (anonymous):

For problems 11–12, find the distance between the two points. Round to the nearest tenth if necessary. 11. (8, 8), (12, 11) (1 point) 7 5 25 28 12. (–2, –6), (3, 9) (1 point) 15.8 250 3.2 20

OpenStudy (callisto):

Use distance formula... \[distance = \sqrt{(x_1 - x_2)^2 +(y_1 - y_2)^2}\]

OpenStudy (callisto):

For example, for 11. \[distance = \sqrt{(8-12)^2+(8-11)^2} = \sqrt{(-4)^2 + (-3)^2}\]\[ = \sqrt{25} =5\] Can you try 12 yourself?

OpenStudy (anonymous):

Distance between two points is calculated by \[distance = \sqrt{(x2-x1)^{^{2}} + (y2-y1)^{2}}\] You can assume either of the points to be x2, x1, y2 and y1 11) \[\sqrt{(12-8)^{2} + (11-8)^{2}}\] \[\sqrt{4^{2} + 3^{2}}\] \[\sqrt{25}\] = 5 answer 11) \[\sqrt{(-2-3)^{2} + (-6-9)^{2}}\] \[\sqrt{-5^{2} + -15^{2}}\] \[\sqrt{250}\] = 15.8 answer